### v1

# Trajectory First: A Curriculum for Discovering Diverse Policies

Abstract
Being able to solve a task in diverse ways makes agents more robust to task variations and less prone to local optima. In this context, constrained diversity optimization has emerged as a powerful reinforcement learning (RL) framework to train a diverse set of agents in parallel. However, existing constrained-diversity RL methods often under-explore in complex tasks such as robotic manipulation, leading to a lack in policy diversity. To improve diversity optimization in RL, we therefore propose a curriculum that first explores at the trajectory level before learning step-based policies. In our empirical evaluation, we provide novel insights into the shortcoming of skill-based diversity optimization, and demonstrate empirically that our curriculum improves the diversity of the learned skills.
1 Introduction
Reinforcement Learning (RL) has driven breakthroughs in robot locomotion [1], game-playing [2, 3], and foundation-model finetuning [4]. While effective, most RL methods assume a unimodal action distribution and produce only a single policy. In contrast, humans and animals routinely solve the same task using multiple qualitatively different strategies. Such variability is also desirable in RL, as it has been shown that human strategy diversity increases solution quality and robustness [5, 6]. Therefore, this work considers the discovery of a policy set that maximizes the reward in diverse ways.
A number of previous works have investigated this problem from various perspectives. Notably, the fields of Novelty Search (NS) and Quality-Diversity (QD) have proposed a multitude of algorithms which populate an archive of solutions based on their novelty and performance [7, 8, 9, 10]. Further, gradient-based RL approaches define intrinsic diversity rewards that they combine with extrinsic task rewards using Lagrange multipliers [11], bandits [12], or linear combinations [13, 14, 15].
While effective, the above approaches are not without shortcomings. QD may produce exceptional results, but often at the cost of sample efficiency and manual feature design. Gradient-based diversity or entropy bonuses in RL may still collapse to a few modes and remain under-evaluated in challenging contact-rich tasks [16]. Since skill learning and diversity or entropy maximization are common exploration objectives (see Zheng et al. [17] for an overview), this observation may appear counterintuitive at first. Yet, we corroborate these findings in our work, demonstrating that even simple obstacle avoidance tasks may be substantially challenging for diversity optimization in RL (Fig. 2). Further, we highlight the necessity for exploration strategies tailored for diversity optimization, since we find that out-of-the-box exploration methods do well in improving task performance, but do not guarantee policy diversity.
To overcome these shortcomings, we propose a new and simple curriculum for learning diverse policies with off-policy and model-free RL (Section 5). We motivate our approach by the observation that skill-based RL methods commonly explore by optimizing neural-network parameters which can be inefficient. In contrast, we introduce a trajectory-first curriculum that (i) uses an evolutionary search over open-loop action sequences to uncover a diverse set of high-reward behaviors, and (ii) distills these behaviors into distinct, off-policy, model-free policies. While prior work proposed similar formulations that first solve exploration and then learning [18, 19], we do not rely on human demonstrations and propose an evolutionary approach to maximize entropy by optimizing trajectory parameters for novelty instead of neural network parameters. Based on this algorithm, we empirically highlight shortcomings of existing diversity optimization methods in Section 5 and illustrate how the proposed curriculum enables learning diverse sets of robot manipulation policies.
In short, we make three contributions with this work. First, we propose a novel curriculum for diversity optimization under extrinsic task rewards (see Fig. 1). Second, we introduce Constrained Novelty Search (CNS) to discover diverse trajectories in the first step of this curriculum (Section 3.1). Finally, we investigate how diversity can be maintained during online training of control policies from this data (Section 3.2).
Figure 1: Overview of the proposed diversity curriculum. First, an evolution strategy explores in trajectory space to maximize novelty of trajectory parameters $\omega\in\Omega$ under performance constraints. Then, this data is used to warmstart the online training of multiple RL agents $\pi\in\Pi$ to solve the same optimization problem in policy space.
2 Preliminaries
Markov Decision Process:
[…]
Constrained Diversity Optimization:
While earlier works used scalars to balance diversity and task rewards, Zahavy et al. [11] introduced the following constrained MDP formulation:
$\max_{\Pi^{n}}\leavevmode\nobreak\ \text{Diversity}(\Pi^{n})\qquad\leavevmode\nobreak\ \leavevmode\nobreak\ \text{s.t.}\leavevmode\nobreak\ \leavevmode\nobreak\ J(\pi)\geq\alpha v^{*},\quad\forall\pi\in\Pi^{n}$ (1)
where $\Pi^{n}$ is the current set of policies, $v^{*}\triangleq\max_{\pi\in\Pi}J(\pi)$ is the value of the optimal policy and $\alpha\in[0,1)$ is a hyperparameter controlling the optimality constraint. This constrained optimization problem can then be solved efficiently using Lagrange multipliers that are tuned using dual ascent [22, 23]. Similar to subsequent work [17], we adopt the same problem formulation in this work. To quantify diversity, we will measure the distance to the nearest neighbor, which shall be maximized:
$\text{Diversity}(\Pi^{n})\coloneq\frac{1}{n}\sum_{i=1}^{n}\min_{\pi_{j}\neq\pi_{i}}\lVert\operatorname*{\mathbb{E}}_{s\sim\rho_{\pi_{i}}}[\phi(s)]-\operatorname*{\mathbb{E}}_{s\sim\rho_{\pi_{j}}}[\phi(s)]\rVert^{2},$ (2)
[…]
Novelty Search:
Most novelty-based approaches to skill learning maximize the entropy of the policy set by using the entropy of the current policy set as an intrinsic reward [10, 8, 27]. To quantify entropy, the particle-based entropy estimator by Singh et al. [28] is commonly employed which estimates the sparsity of the distribution based on the distance between the datapoints $\{x_{i}\}_{i=1}^{n}$ and their $k$-nearest neighbor: ${\cal H}_{\text{particle}}(X)\propto\sum_{x_{i}\in X}\log\lVert x_{i}-x_{i}^{(k)}\rVert$. In practice, we choose $k=1$. In other words, we measure the particle-based entropy as distance to the nearest neighbor, and add a constant $c=1$ for numerical stability:
${\cal H}_{\text{particle}}(X)\coloneq\sum_{x_{i}\in X}\log\left(c+\min_{x_{j}\neq x_{i}}\lVert x_{i}-x_{j}\rVert\right).$ (3)
3 Evolutionary Exploration for Diverse Policy Discovery
Solving the problem in Eq. 1 requires an initialization that is sufficiently diverse to prevent the diversity optimization from only occurring locally. Since the policies are initialized randomly at the beginning, we find that the discovered policy set rarely covers the task space sufficiently and instead focuses on a subset of greedy solutions. While shaping the extrinsic reward is an option to encourage diverse interactions between agent and environment, tuning such reward functions is a tedious task. We propose an alternative approach to this, which $(i)$ uses an evolution strategy (ES) to explore in the space of trajectories before $(ii)$ learning a diverse set of policies from this data. We provide an intuition for this approach in Fig. 1 and describe both stages of our curriculum in the following. For a formal algorithmic description of the method, we refer to Appendix A.
3.1 Constrained Novelty Search for Spline-based Exploration
The first stage of our curriculum directly optimizes agent trajectories $\tau\in{\mathbb{R}}^{T\times u}$, where $T$ denotes the number of timesteps and $u$ the robot’s degrees of freedom. We represent a trajectory as a B-spline [29] parameterized by a control point matrix $\omega\in{\mathbb{R}}^{m\times u}$. Following the constrained diversity optimization objective in Eq. 1, we optimize a set of trajectory parameters $\Omega=\{\omega_{i}\}_{i=1}^{n}$ such that the resulting trajectories $\tau(\omega)$ are as diverse as possible under the constraint of near-optimality:
$\max_{\Omega^{n}}\leavevmode\nobreak\ \text{Diversity}(\Omega^{n})\qquad\leavevmode\nobreak\ \leavevmode\nobreak\ \text{s.t.}\leavevmode\nobreak\ \leavevmode\nobreak\ J(\tau(\omega))\geq\alpha v^{*},\quad\forall\omega\in\Omega^{n},$ (4)
where $J(\tau(\omega_{i}))=\textstyle\sum_{t}r_{\text{ext}}(s^{i}_{t},a^{i}_{t})$, and $v^{*}=\max_{\omega}J(\tau(\omega))$. Since the extrinsic task reward function $r_{\text{ext}}$ is generally non-differentiable we optimize this problem using a multi-population evolution strategy (ES). Following Zahavy et al. [11] we solve the dual optimization using gradient ascent with bounded Lagrange multipliers $\{\lambda_{i}\}_{i=1}^{n}$, which yields the following trajectory reward that we evaluate on the trajectory $\tau_{i}=\tau(\omega_{i})$:
$r(\tau_{i})$ $=(1-\sigma(\lambda_{i}))\leavevmode\nobreak\ r_{\text{int}}(\tau_{i})+\sigma(\lambda_{i})\leavevmode\nobreak\ r_{\text{ext}}(\tau_{i})$ (5)
$=\sum_{s_{t}\in\tau_{i}}(1-\sigma(\lambda_{i}))\leavevmode\nobreak\ {\cal H}_{\text{particle}}(\phi(s_{t}))+\sigma(\lambda_{i})\leavevmode\nobreak\ r_{\text{ext}}(s_{t})$ (6)
$=\sum_{s^{i}_{t}\in\tau_{i}}\left[(1-\sigma(\lambda_{i}))\log\left(1+\min_{\tau(\omega_{j})\in\Omega}\lVert\phi(s_{t}^{i})-\phi(s^{j}_{t})\rVert_{2}\right)+\sigma(\lambda_{i})\leavevmode\nobreak\ r_{\text{ext}}(s_{t})\right],$ (7)
where $\phi$ is a feature extraction function that projects the states to a lower dimension, $\lambda_{i}$ is the $i$-th Lagrange multiplier, and $\sigma$ denotes the sigmoid function $\sigma(x)=1/(1+\exp(-x))$. We note that Eq. 7 is a generalized version of the novelty search objective from Conti et al. [10], but using population-level Lagrange multipliers instead of a single heuristically selected scalar. Using Lagrange multipliers permits not only to consider different weights for different populations, but also a dynamic adaptation of these weightings. We denote this objective and its optimization as Constrained Novelty Search (CNS) in the following. While Zahavy et al. [11] derive the analytical gradient of Eq. 2 for their intrinsic reward, CNS approximates this gradient by stochastic sampling from an ES, which approximates natural gradient steps on the novelty objective [30, 31, 32].
[…]
Finally, we note again that we optimize in ${\mathbb{R}}^{m\times u}$ instead of ${\mathbb{R}}^{d}$ at this step where $d$ is the dimension of policy parameters. Since typically $m\times u\ll d$, we can optimize Eq. 7 with the CMA-ES [32], which is empirically more sample efficient in moderately high dimensional parameter spaces than using isotropic search distributions [35, 10] .
3.2 Efficient Online Diversity Optimization from Prior Data
Given a diverse dataset ${\cal D}=\{(\tau_{i},\,z_{i},\,r_{\mathrm{ext}}(\tau_{i}))\}_{i=1}^{\ell}$ with skill labels $z_{i}\in\{1,\dots,n\}$ from CNS, our second stage learns $n$ reactive, skill-conditioned policies that preserve diversity while satisfying the near-optimality constraint. We build on the off-policy Domino framework [11], and augment it with three key modifications inspired by efficient offline-to-online RL [36].
Following Zahavy et al. [11], we use the gradient of the diversity objective as intrinsic reward for policy optimization, that is
$r_{\text{int}}(s_{t},a_{t}\mid z)=\phi(s_{t})^{\top}\left(\bar{\phi}_{z}-\bar{\phi}_{j}\right),\quad\leavevmode\nobreak\ \leavevmode\nobreak\ \text{s.t.}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \bar{\phi}_{j}=\operatorname*{arg\,min}_{j\neq z}\lVert\bar{\phi}_{z}-\bar{\phi}_{j}\rVert^{2}_{2},$ (8)
where $\bar{\phi}_{z}=\operatorname*{\mathbb{E}}_{s\sim\rho_{\pi_{z}}}[\phi(s)]$ are the expected features per skill. For consistency, the intrinsic term measures novelty in the same projection space $\phi$ used in CNS. We balance extrinsic and intrinsic rewards as in Eq. 7 using bounded Lagrange multipliers that we update via dual ascent to enforce the near-optimality constraint. Similar to Domino, we track a running estimate of the expected features per skill, which we initialize with the population means from CNS.
To efficiently incorporate the CNS data, we use symmetric sampling [36, 37, 38], which means that each batch is composed of equal parts of online and offline transitions. Unlike prior work, we do not use the full data, however, since we found that trajectories from early iterations of CNS might fail to make meaningful interactions with the environment. So we only keep those trajectories that are nearly optimal following Eq. 1. To also exploit suboptimal data, we use a relaxed near-optimality criterion of $\alpha/4$ and estimate $v^{*}=\max_{\tau\in{\cal D}}J(\tau)$. Since this procedure rejects unequal numbers of trajectories per-skill, we adapt the symmetric sampling by balancing offline and online data equally per-skill. Further, we use a high number of learning steps per environment step (update-to-data ratio, UTD) to learn from the diverse CNS data as efficiently as possible. This allows to propagate exploration data quickly through the network, but requires extensive regularization to prevent overfitting. While multiple remedies to this issue are known, we follow Ball et al. [36] in using random ensemble distillation [39], and observation and layer normalization.
4 Related Work
Diversity-Driven Policy Discovery.
Various methods to search diverse policies have been proposed. Quality-Diversity (QD) and evolutionary methods search in a gradient-free manner, populating archives of high-performing, behaviorally distinct solutions [40, 41] or co-optimizing fitness and novelty across populations [12, 10, 42]. In principle, these approaches could be used in the first step of the proposed curriculum. However, most prior work focuses on optimizing in policy parameter space before distilling policies [43, 44, 45], a less effective process as we find in this work. Recently, gradient-based RL has been reformulated to discover multiple policies via intrinsic diversity bonuses. Notably, Eysenbach et al. [26] maximize mutual information between skills and states, but focus on unsupervised skill discovery, while we target task-driven diversity in this work. For environments with extrinsic rewards, prior methods either learn policies sequentially [46, 14, 47, 48], or in parallel for greater efficiency [11, 15, 48]. Further, the methods differ in the employed optimization paradigm. While earlier works balance extrinsic and intrinsic rewards with fixed scalars [14, 49], more recent works proposed adaptive weighting schemes [12, 13]. In particular, Zahavy et al. [11] proposed constrained optimization with Lagrange multipliers, which we adopt in this paper.
Exploration in RL.
Exploration is a fundamental aspect of RL, enabling agents to effectively sample the environment, avoid premature convergence to suboptimal policies, and enhance both learning performance and generalization. Accordingly, numerous exploration strategies have been proposed in the literature [50]. A common strategy perturbs the agent’s actions – often via Gaussian or temporally correlated noise processes [51, 52]. Another line of work introduces parameter noise, where noise is applied directly to the agent’s parameters rather than to its actions [53, 54]. Beyond pure noise, intrinsic-reward methods augment the extrinsic task reward with bonuses for novelty. These approaches include techniques based on knowledge-based exploration, which maximizes prediction error [55], competence-based exploration [56, 26, 57, 17] and data-based exploration, which maximizes entropy [27]. While the above intrinsic exploration objectives or parameter noise explore in the parameter space, we propose to perform intrinsically motivated exploration in the much lower-dimensional trajectory space. By optimizing a novelty objective on trajectories using an ES, we can explore over entire behaviors, much like temporally correlated action noise, but with a self-optimizing noise distribution. Finally, unlike methods that rely on expert demonstrations to guide exploration [19, 58], our approach does not require prior knowledge or external supervision.
RL Finetuning from Datasets.
Offline RL addresses the issue of data inefficiency inherent in online RL by training solely on a fixed dataset of past interactions, but it often suffers from sub-optimality due to limited or biased data [59]. To overcome these problems, the paradigm of offline-to-online RL has been proposed, where a policy is trained on offline and online data in conjunction [60, 61, 62, 63, 59, 64, 65]. The approaches broadly fall into two categories based on whether the dataset from the replay buffer is discarded after a pretraining phase, or whether it is retained [61]. Pretrain-and-discard methods first pretrain a policy or critic on the dataset and then use the same networks for finetuning without retaining any prior data [47, 66, 67]. Data retention methods keep the dataset in the replay buffer for at least a fraction of the training procedure. Many methods that fall into this category first employ an offline pre-training phase but also retain the data during online training to prevent catastrophic forgetting [60, 59, 68]. Other works instead directly train a policy online, mixing samples from the prior dataset and current rollouts [36, 69, 62, 37, 19]. We adopt this mixing strategy for our CNS dataset finetuning but observe that none of these approaches explicitly address the preservation or enhancement of policy diversity during online adaptation. In this work, we close that gap by investigating how offline–online data selection and mixing influence the retention and amplification of diverse behaviors.
5 Experiments
(a) Cylinder Avoidance Task
(b) No Exploration (Domino)
(c) Online Exploration ($\text{Domino}+\epsilon\sim{\cal N}$)
(d) Curriculum Exploration (Ours)
Figure 2: Qualitative results for the cylinder avoidance task. Each plot depicts $xy$ trajectories around black obstacles. Each of the 10 skills is plotted in a separate color. Curriculum exploration is the only method that finds all collision free paths through the maze.
(a) Cube Pushing Task
(b) No Exploration
(c) Online Exploration
(d) Curriculum Exploration (Ours)
Figure 3: Qualitative results for the cube pushing task. Each plot depicts the rollouts of 4 skills that were trained on this task. Curriculum exploration is the only method that maximizes diversity by pushing the block to four different corners of the table.
Our experiments are designed to provide insights into the shortcomings of naively applying diversity constraints and how these can be mitigated. As such, we aim to answer the following questions: (a) Does constrained diversity optimization sufficiently explore the environment? (b) Does a curriculum with constrained novelty search increase policy diversity in skill learning? (c) Do we need diversity objectives and performance constraints?
Environments. We conduct our experiments focusing on environments from robotics. We use two environments for evaluation: (a) a cylinder avoidance task where a rod attached to a robot gripper must be navigated through a maze without collisions (see Fig. 2). The agent is rewarded for minimizing the distance to the goal line and penalized for touching obstacles. This task enables a rich set of behaviors since many paths through the maze are possible. To avoid the trivial solution of moving over the top of all obstacles, we fix the $z$-position of the rod and use $xy$-endeffector position control. We train 10 different skills for this task. (b) Second, we use a cube pushing task in which a 3d-pointmass robot is tasked to push a cube as far away from the center of a table as possible. This task enables a rich set of behaviors since there are no contact encouraging terms in the reward, so any object manipulation is conceivable. We train 4 skills for this environment. For full environment details, we refer to Appendix B.
Baselines. We compare our method to two baselines: First, we use plain diversity optimization with optimality constraints, without additional exploration, to investigate how well these methods explore. Second, we use Gaussian action noise $\epsilon\sim{\cal N}(0,I)$ as a non-diversity-specific exploration method. We implement the baseline diversity optimization as described in Domino [11], and base our code on the public implementation thereof [70]. For both baselines, we do not estimate successor features and use the ground truth observations as state features to guarantee maximum possible performance [11, 71]. As stated above, all code is based on SAC [21], and uses observation normalization, critic ensembling, and layer norm. For further implementation details, we again refer to Appendix B.
Q1. Does constrained diversity optimization sufficiently explore the environment?
To answer this question, we look at the performance of running diversity optimization without any additional exploration mechanism. We observe in Fig. 2 that the diversity optimization algorithm Domino underexplores the domain, which leads to little behavioral diversity. While the method successfully solves the problem, only a few behavioral modes are found. For the cube task, we see that multiple distinct behaviors can be found, but some skills overlap (Fig. 3). This highlights that learning a set of diverse skills requires the discovery of a sufficient number of behaviors in the first place. Further, our quantitative results (Fig. 4) highlight that online exploration only partly resolves this issue, as the overall diversity of the solutions is still limited, which we can also see in Fig. 2.
Q2. Does a curriculum with constrained novelty search increase policy diversity?
To answer this question, we consider the results of training agents using the proposed curriculum. As we can see in Fig. 2, using the curriculum enables to discover almost all paths though the maze. Similarly, in Fig. 3, the curriculum exploration is the only method that finds pushes along four different directions of the table. All other methods tend to explore more subtle variations such as the final orientation of the cube once it is pushed off the table. These qualitative findings are corroborated quantitatively in Fig. 4, which shows that the proposed curriculum clearly produces the most diverse policies.
(a) Cylinder Avoidance Task
(b) Cube Pushing Task
Figure 4: Quantitative evaluation. Curriculum exploration leads to higher policy diversity at high performance. We report interquartile mean (IQM) and 95% confidence intervals across 5 seeds.
Q3. What is the value of diversity objectives?
This question aims to investigate whether the diversity objective during RL is needed given diverse data from the first stage of the curriculum. To answer this question, we run constraint-free population-based training [72, PBT] based on the CNS data, but only maximizing policy return. Our results in Fig. 5a show that using the diversity objective indeed increases the policy diversity. At the same time, we observe a slight decrease in task performance when optimizing for diversity, which we explain by the objective from Eq. 1, which permits a certain amount of slack $\alpha$. We believe that these observations provide valuable insights into RL finetuning for diversity. As stated above, prior finetuning literature neglected diversity, and we believe that the presented recipe closes this gap. Further, we investigate whether the inclusion of the Lagrange multiplier in the CNS formulation improves the evolutionary trajectory optimization in comparison to the scalarized novelty search formulation from previous work [10]. Fig. 5b displays the population entropy and population reward across novelty search iterations. We observe that while there are no differences in rewards, the population entropy is higher for the proposed CNS objective. This demonstrates that the generalized formulation of novelty search may be a valuable tool for exploration and discovery in the future.
(a) RL Diversity Objective Ablation
(b) Lagrangian Ablation
Figure 5: Ablation study on cube pushing task. We compare performances for RL with diverse data from CNS, but no diversity objective during training (PBT) against our proposed approach. Further, we ablate the usage of Lagrange multipliers.
6 Conclusion & Limitations
We introduced a two-stage, trajectory-first curriculum that can discovering diverse skills in challenging robotic domains. In the first phase, we use a constrained novelty search evolution strategy to explore trajectories. In the second stage, we train a set of diverse reactive control policies given the CNS data. We have shown in our experiments that: (a) naïvely applying constrained diversity objectives in policy space leads to under-exploration and thus fails to discover truly diverse skills (Fig. 2). (b) By first exploring diverse trajectories using constrained novelty search, the diversity optimization can be improved (Fig. 4). (c) Using proper performance constraints during evolutionary novelty search improves sample set entropy over scalar formulations (Fig. 5a). (d) Using diverse data for policy training alone does not guarantee truly diverse policies, but diversity objectives are still required to maintain full diversity (Fig. 5b).
Despite these advances, our approach is not without limitations. We observe a high variance in task returns (Fig. 4), which we hope to alleviate in the future. For instance, it may be possible to improve the data-based finetuning part of our pipeline, by simply increasing the number of seeds in our evaluation. Moreover, despite providing valuable insights, we aim to further understand how exploration can improve diversity optimization. We plan to compare to additional baselines such as data-based exploration methods [27, 55]. Finally, our method introduces new hyperparameters to the diversity optimization. In the future, we aim to conduct sensitivity analyses to gain a more profound understanding of the approach’s hyperparameters.
Acknowledgments and Disclosure of Funding
This research was funded by the Amazon Development Center Germany GmbH.
Appendix A Algorithm
Algorithm 1 Curriculum for discovering diverse policies.
Input: Environment $env$, Optimality ratio $\alpha$, Num. skills $n$, Learning rates $\kappa_{\lambda},\kappa_{\xi},\kappa_{V},\kappa_{\pi}$, Init. SAC temperature $\xi$
1:// 1. Constrained Novelty Search
2:Initialize population parameters $\omega_{i}$ for skills $i=1,\dots,n$
3:${\cal D}_{cns}\leftarrow\{\}$
4:for iteration $t=1,\dots,T$ do
5:   for Population $i=1\dots n$ do
6:     $\{\tau^{i}_{1},\dots\tau^{i}_{m}\}\leftarrow\texttt{env.rollout}(\omega_{i})$
7:     Update $\omega_{i}$ given $(1-\sigma(\lambda^{cns}_{i}))\leavevmode\nobreak\ r_{\text{int}}(\{\tau^{i}_{1},\dots\tau^{i}_{m}\})+\sigma(\lambda^{cns})\leavevmode\nobreak\ r_{\text{ext}}(\{\tau^{i}_{1},\dots\tau^{i}_{m}\})$ (Eq. 4)
8:     $v_{i}\leftarrow v_{i}+\kappa_{\lambda}\operatorname*{\mathbb{E}}[r_{ext}(\{\tau^{i}_{1},\dots\tau^{i}_{m}\})]$ (Estimate population values)
9:     if t % LAMBDADELAY then
10:       $\lambda^{cns}_{i}\leftarrow\lambda^{cns}_{i}-\alpha_{\lambda}\nabla(\lambda^{cns}(v_{i}-v^{*}\alpha))$
11:     end if
12:     ${\cal D}_{cns}\leftarrow{\cal D}_{cns}\cup\{\tau^{i}_{1},\dots\tau^{i}_{m}\}$
13:   end for
14:end for
15:
16:// 2. Constrained RL Diversity Optimization
17:for iteration $t=1,\dots,I$ do
18:   // Environment steps
19:   $z\sim p(z)$
20:   $a_{t}\sim\pi(a_{t}\mid s_{t},z)$
21:   $s_{t+1}\sim p(s_{t+1}\mid s_{t},a_{t},z)$ (Step environment)
22:   ${\cal D}_{rl}\leftarrow{\cal D}_{rl}\cup\{(s_{t},a_{t},r(s_{t},a_{t}),\phi(s_{t},a_{t}),s_{t+1},z)\}$
23:   // Training steps
24:   ${\cal D}_{batch}\leftarrow{\cal D}_{1}\cup{\cal D}_{2}\text{ with }{\cal D}_{1}\sim{\cal D}_{cns},{\cal D}_{2}\sim{\cal D}_{rl}$ (Symmetric sampling)
25:   $\xi\leftarrow\xi-\kappa_{\xi}\,\nabla J_{\xi}(\xi)$ (Update SAC temperature)
26:   $\lambda^{rl}\leftarrow\lambda^{rl}-\kappa_{\lambda}\,\nabla J_{\lambda}(\lambda^{rl})$ (Dual ascent on Eq. 1)
27:   $\theta_{V}\leftarrow\theta_{V}-\kappa\,\nabla J_{V}(\theta_{V})$ (Update critic)
28:   $\theta_{\pi}\leftarrow\theta_{\pi}+\kappa_{\pi}\,\nabla J_{\pi}(\theta_{\pi})$ (Update policy)
29:end for
Appendix B Experimental Details
Each experiment is repeated across 5 different seeds. Where applicable, we report the interquartile mean (IQM) across all 5 runs and bootstrapped 95% confidence intervals in our plots. In the following we provide details about the environments and implementation that we used in this work. All the experiments are performed on an internal cluster with eight NVIDIA A40 GPUs. We evaluate our method on two robotics tasks, which are displayed in Fig. 6. Both environments requires reasoning over objects in the scene, which is generally challenging.
B.1 Environments
Figure 6: Considered environments. Left: cylinder avoidance task. Right: cube pushing.
Cylinder Avoidance Task.
In this task, a robot must successfully navigate a rod that is attached to its gripper around 6 cylinder obstacles without knocking them over. The agent is rewarded for minimizing the distance to the goal line and penalized for touching obstacles. To avoid the trivial solution of moving over the top of all obstacles, we fix the $z$-position of the rod and use $xy$-endeffector position control. Therefore, the action space is $a\in[-1,1]^{2}$, while the observation space are position and velocity information, i.e., $s\in{\mathbb{R}}^{4}$. We train 10 different skills for this task. The reward function that we use is the following:
$r(s,a)=s_{x}-target_{x}-\beta\cdot\mathds{1}_{collision}(s)+x_{max},$
where $s_{x}$ denotes the $x$ position of the rod, $target_{x}$ is the coordinate of the goal line, while $\mathds{1}_{collision}(s)$ is a collision checking function. Further $x_{max}$ is the maximum $x$ coordinate that is admissible, which we use as offset to guarantee position rewards. We additionally clip the rewards to be in $[0,2]$ to improve training stability. We choose $\beta=20$ for our experiments. To further simplify the task, we bound the $xy$-positions to $[-4.5,4.5]^{2}$, which we implement by clipping.
Cube Pushing Task
In this task, a 3d-pointmass robot is tasked to push a cube as far away from the center of a table as possible. The agent is only rewarded for maximizing the distance between the cube’s center of mass and that of the table. This task is challenging because the reward is spatially sparse as the robot must detect the cube by exploring the environment. The action space is $a\in[-1,1]^{3}$, while the observation space are position and velocity information of robot and cube, i.e., $s\in{\mathbb{R}}^{19}$. We train 4 skills for this environment. The reward function is defined as follows:
$r(s,a)=\dot{q}_{cube}+\beta(\lVert xyz_{cube}-xyz_{table}\rVert).$
Here $\lVert xyz_{cube}-xyz_{table}\rVert$ denotes the position difference between cube and table center of mass. We approximate $\dot{q}$ by first order finite differences as $\dot{q}(x)\approx x_{t+1}-x_{t}$, which we find to produce better training results than using the velocities from the Mujoco simulator that we use [73]. We choose $\beta=1/2$.
B.2 Implementation
We implement all algorithms in JAX [74]. We implement Constrained Novelty Search based on the CMA-ES implementation from evosax [75]. The code for all RL agents is based on Domino [11], the public implementation thereof Grillotti et al. [70], and the STOIX ecosystem [76]. We follow Zahavy et al. [11] in the choices of all hyperparameters for Domino with exceptions detailed below. For all baselines, we use ground truth observations as state features, since they are low-dimensional and should thus yield the best performances [11, 71]. For CNS, we use the aforementioned random projections since they are an elementary part of the method. For a full list of hyperparameters, we refer to Table 1.
Initialization
Since we initialize Domino from prior data, we adapt the initialization. The running estimates of the state features are not initialized with $\phi^{avg}=\bar{1}/f$ for features in ${\mathbb{R}}^{f}$. Instead, we use the mean of the maximum likelihood solutions from the CNS. In other words, for each CMA-ES population that we run, we select the resulting trajectory parameters, roll out an additional trajectory from them and use the expected features over this trajectory as initial estimate of the features per skill. We find that this initialization is provides better results than the uniform initialization from Domino. For the values however, we follow Domino in using a zero initialization for all skills instead of using the expected reward from the final parameter rollouts. This is because such an initialization would overestimate the capacities of the current policies and thus only optimize diversity from the very beginning of training following Eq. 1.
Network Architectures
As stated in Section 3.1 we follow design choices from prior work in using layer normalization and observation normalization. We use the same architectures for actor and critic networks, however we perform critic ensembling for regularization and use separate heads for extrinsic and intrinsic values. Our MLP backbone resembles that of Lee et al. [77], but we replace the layer norm in between block by a dynamic tanh [78]. All networks are optimized using Adam [79].
Constrained Novelty Search.
We implement CNS based on the CMA-ES implementation in evosax. Before combining extrinsic and intrinsic fitness, we normalize both values within each subpopulation. Further, we use simple gradient descent to update the Lagrange multipliers. For higher optimization stability, we only update these parameters every iteration, but then perform 200 steps of gradient descent. To prevent gradient saturation due to the usage of sigmoids on the Lagrange multipliers, we bound them to make sure that they remain in a reasonable range. Similar to Domino, we also fix the first Lagrange multiplier to $1$, so we can estimate $v^{*}$ based on this population. In practice, we found it more stable, however, to choose $v^{*}=\max_{\omega\in\Omega}v(w)$. Note that we follow this choice only during CNS.
Appendix C Additional Results
Sample Efficiency
We now present additional experimental results evaluating the sample efficiency of our method. To this end, we plot the learning curves with respect to the extrinsic task in Fig. 7. We can see that on the simpler task of cylinder avoidance, the final task performance is similar, but is less sample-efficient. In the cube task, however, in which it is more challenging to explore, there is no difference in sample efficiency between the methods. This highlights the potential of the proposed curriculum.
(a) Cylinder Avoidance Task
(b) Cube Push Task
Figure 7: Sample efficiency comparison. While the ES-free method achieves higher performance earlier on the simpler cylinder avoidance task, the sample efficiency on the cube push task is similar.
Full Results
For completeness, we report the results of every seed for the obstacle avoidance task in Fig. 8. We see that our method learns the most diverse policy set for each seed.
Seed01234
No Exploration
Online Exploration
Curriculum Exploration
CNS Data
Figure 8: Complete results for cylinder avoidance task. Each column is a different seed and each row a different method. The first row depicts running Domino without additional exploration. The second row uses action noise for improved online exploration. The third row shows our approach, which learns the most diverse trajectory set for each seed. In the bottom row, we depict the data which we obtain from the CNS optimization. For visibility reasons, we subsample the data by a factor of 10.
Cylinder Avoidance Cube Pushing
Environment Details
   Observation size 4 19
   Action size 2 3
   Episode length 100 200
   Num. env. steps 800 000 2 500 000
   Num. skills 10 4
   Optimality ratio $\alpha$ 0.75 0.6
RL Parameters
   Update-to-data ratio 10 10
   Discount 0.95 0.99
   Batch size 32 32
   ${\cal H}_{target}$ $\dim{\cal A}$ $\dim{\cal A}$
   Critic hidden depth 2 2
   Critic hidden size 256 256
   Actor hidden depth 2 2
   Actor hidden size 256 256
   Learn. Rate Critic 3e-4 3e-4
   Learn. Rate Actor 3e-4 3e-4
   Learn. Rate Temperature 3e-4 3e-4
   Learn. Rate Lagrange 1e-3 1e-3
   Optimizer Adam Adam
   Polyak weight 0.005 0.005
   Num. critics 8 8
   Critic subset size 2 2
CNS Parameters
   Num. iterations 100 100
   Subpopulation size 4 8
   Elite ratio 0.4 0.4
   Init. $\sigma$ 0.8 0.8
   Random feature dim. 2 4
   Lagrange range [-2, 2] [-2, 0]
   Num. spline controls 10 8
Table 1: Full Hyperparameter Overview